### v1

# Resonant Solitary States in Complex Networks

Abstract
Partially synchronized solitary states occur frequently when a synchronous system of networked oscillators is perturbed locally. Remarkably, several asymptotic states of different frequencies can coexist at the same node. Here, we uncover the mechanism underlying this multistability: The resonant back-reaction of the network’s modes on the solitary can lead to a large energy transfer between them, and the solitary’s frequency adaptation can tune the system to this resonance. We provide a full analytic analysis of this mechanism, and show which network structures enable its presence.
Oscillator systems are ubiquitous in the description of natural and human-made systems. Important extended systems as diverse as the heart, the brain, firefly populations, chemical reactions, and power grids, can be modelled as networks of interconnected oscillators Strogatz (2000); Acebrón et al. (2000, 2005); Pecora et al. (2014); Rodrigues et al. (2016). One of the most fundamental common aspects of the collective behavior of such systems is synchronization Pikovsky et al. (2001). In some, such as the power grid, it is essential for their proper functioning, in others, such as the brain, it can indicate severe dysfunction. The tendency to synchronize, and thus the function of the system, is strongly influenced by the underlying network’s topology.
The paradigmatic models for understanding this relationship between structure and function in coupled oscillator networks are the Kuramoto model Kuramoto (1984) and its variants. These feature extremely rich collective behavior, such as chimera states, frequency clusters, isolated desynchronization, and spatial chaos Strogatz (2000); Acebrón et al. (2000, 2005); Pecora et al. (2014); Rodrigues et al. (2016). The Kuramoto model with inertia Tanaka et al. (1997a, b) has been developed independently to study synchronization properties in power grids Bergen and Hill (1981); Witthaut et al. (2022) and biological systems Ermentrout (1991).
The transition from decoupling to synchrony in inertial Kuramoto networks is characterized by spatial chaos, a form of extreme multistability Omelchenko et al. (2011). This regime typically has several states with similarly sized basins of attraction Gelbrecht et al. (2020); Halekotte et al. (2021). Among the most prominent are frequency clusters Berner et al. (2021); Olmi et al. (2014), the simplest of which are solitary states. In a solitary state, only one or a few oscillators start to rotate at their own frequencies, while the rest of the network remains synchronized Maistrenko et al. (2014); Jaros et al. (2015). Solitary states are especially prominent for the important situation of localized large perturbations Menck et al. (2014); Hellmann et al. (2020); Jaros et al. (2018), and they are particularly likely for perturbations at leaf nodes Menck et al. (2014); Nitzbon et al. (2017). Localized perturbations are important in systems like power grids, where single component failures are common and can lead to desynchronization and blackouts UCTE .
Interestingly, numerical studies have revealed that solitary nodes can exist not only at the natural frequency of the oscillators, but also at shifted frequencies in networks with losses Hellmann et al. (2020) and at intermediate frequencies Nitzbon et al. (2017). In addition, the presence of multistability in these solitary nodes has been identified, allowing for the coexistence of both intermediate and natural solitary states Nitzbon et al. (2017); Halekotte et al. (2021) at the same node.
Previous work has discussed the role of a networks’ topology in understanding its overall response to perturbations and the emergence of various stable states Rohden et al. (2012); Menck et al. (2014); Rodrigues et al. (2016); Nitzbon et al. (2017); Zhang et al. (2019). In some contexts, the frequency and stability of the solitary states have been studied in terms of decoupling and entrainment arguments Menck (2014); Hellmann et al. (2020). These explanations do not account for either intermediate frequency solitaries or the coexistence of several solitary states at the same node. In the following, we use averaging theory Guckenheimer and Holmes (2002) and linear response to develop a theory of the resonant coupling between the synchronized cluster and the solitary. We uncover that the described phenomena arise from the cluster’s resonantly excited complex network structure. This leads to a non-zero mean energy flow between cluster and solitary, and a frequency shift (Fig. 1). While works like Gao and Efstathiou (2018); Yue et al. (2020); Munyayev et al. (2022) have previously considered the interaction between individual oscillators and the mean field, such an interaction between collective network modes and individual oscillators has not been previously described.
Figure 1: Frequencies ($\dot{\varphi}_{i}$) and time-averaged energy flow $p_{s}$ between solitary and synchronous component for a complex oscillator network following two single-node frequency perturbations, see supplementary material (SM) for an animation and the network topology (Fig. S2) Sup . The first perturbation brings the system from the synchronized regime into an intermediate solitary state in resonance with the synchronized cluster. The second perturbation causes the transition to a natural solitary state, where the solitary node is effectively decoupled and the oscillations of the synchronized cluster are much smaller. The distinct blue trajectory belongs to the solitary node 75 in Fig. 2.
In the following, we will use the Kuramoto model with inertia. This is given by the coupled second-order equations
$m_{i}\ddot{\varphi}_{i}$ $=P_{i}-\alpha_{i}\dot{\varphi}_{i}-\sum_{j=1}^{N}p_{ij}(\varphi_{i},\varphi_{j}),$ (1)
where $\varphi_{i}$ are the phases of the $N$ oscillators, $P_{i}$ the driving powers, the $m_{i}$ are inertia constants, and the $\alpha_{i}>0$ are damping coefficients. The coupling is given in terms of the weighted coupling matrix $\left\{\kappa_{ij}\right\}$ as $p_{ij}(\varphi_{i},\varphi_{j})=\kappa_{ij}\sin(\varphi_{i}-\varphi_{j})$.
In a synchronous state, $\varphi_{i}^{*}$, $\dot{\varphi}_{i}^{*}=\hat{\omega}$, the phases satisfy $P_{i}-\alpha_{i}\hat{\omega}=\sum_{j=1}^{N}p_{ij}(\varphi_{i}^{*},\varphi_{j}^{*})$. By going to a corotating frame, we can always set $\hat{\omega}=0$, and will assume so from now on. The $p_{ij}(\varphi_{i}^{*},\varphi_{j}^{*})$ can be interpreted as the energy flowing through the network to balance out the driving powers $P_{i}$, (this is the physical interpretation for power grids), and $\alpha_{i}$ determines the restoring force of the frequency. The natural frequencies of the uncoupled oscillators (i.e., when all $\kappa_{ij}=0$) are determined by the condition that driving power and restoring force balance: $\Omega_{i}\coloneqq P_{i}/\alpha_{i}$.
Figure 2 shows the asymptotic frequencies actually observed following single node perturbations in a complex network generated with Schultz et al. (2014) for which $\Omega_{i}=\pm 10$. When perturbations occur at nodes 74, 75 and 84, they induce intermediate frequencies. These nodes are leaves with a highly connected neighbor. Further, intermediate frequencies occur for all such nodes in this network.
Figure 2: Asymptotic solitary frequencies in a synthetic power grid after perturbations. Three solitary states are clearly distinct from the natural frequencies at ±$\pm 10$. They occur at leaf nodes with highly connected neighbors (cf. Nitzbon et al. (2017)) nodes 74, 75 and 84, shown in the insets. Nodes shown as ’$+$’ (’--’) have $P_{i}=1$ $(-1)$. States are counted as desynchronized if there are frequencies whose absolute value exceeds the threshold $\omega_{th}=1$, and as solitary if there is exactly one desynchronized node, cf. Fig. S4. Shown here is the absolute value of $\omega_{s}$, which coincides with $\omega_{s}\cdot\text{sgn}(P_{i})$ for each node $i$. See Fig. S2 for the full network, and Fig. 1 for a solitary state of node 75.
Ansatz and averaging.— We will begin by introducing an ansatz for describing a solitary leaf node $z$ and a synchronized cluster $S=\{i:1\leq i\leq N,i\neq z\}$. For simplicity, we will assume that $\alpha_{i}$ and $m_{i}$ are constant throughout $S$. Node $z$ is connected to $S$ through the intermediary node $k\in S$. The coupling between them is given by $p_{zk}$, which we decompose into a long-term time average $p_{s}=\langle p_{zk}\rangle$ and a mean zero part: $p_{zk}(t)\coloneqq p_{s}+p_{zk}^{\text{osc}}(t)$. Our ansatz is
$\varphi_{i}(t)$ $=\omega_{\text{sync}}t+\vartheta_{i}^{*}+\vartheta_{i}(t)\quad\text{for $i\in S$,}$ (2)
$\varphi_{z}(t)$ $=\omega_{s}t+\vartheta_{k}^{*}+\vartheta_{z}(t).$ (3)
Here, we define $\omega_{\text{sync}}$ and $\omega_{s}$ as the long term average frequency of the synchronized cluster, and the solitary oscillator, respectively. The $\vartheta_{i}^{*}$ are a synchronous state of $S$ given a fixed power injection $p_{s}$ at $k$.
Solitary states are those for which $\vartheta_{i}(t)$ and $\vartheta_{z}(t)$ are bounded and the corotating solitary frequency, defined as $\omega_{c}\coloneqq\omega_{s}-\omega_{\text{sync}}$, is non-zero. It can also be shown by summing over Eq. (1), that $(N-1)\alpha\omega_{\text{sync}}+\alpha_{z}\omega_{s}=0$. This gives us
$\omega_{c}(\omega_{s})=\left(1+\frac{\alpha_{z}}{(N-1)\alpha}\right)\omega_{s}.$ (4)
Now, a central step is to linearize the equations around $\vartheta_{i}=0$. This is justified numerically, as the synchronized oscillators remain relatively close to the synchronous state. To linear order in $\vartheta_{k}$, $p_{zk}$ is
$p_{zk}(t)$ $=\kappa_{zk}\left\{\sin\left[\omega_{c}t+\vartheta_{z}(t)\right]-\vartheta_{k}(t)\cos\left[\omega_{c}t+\vartheta_{z}(t)\right]\right\}.$ (5)
Introducing the vectors $e^{k}_{i}:=\delta_{ik}$, and $\vec{\vartheta}$ with components $\vartheta_{i}$ for all $i\in S$, the system is
$m\ddot{\vec{\vartheta}}=$ $-\alpha\dot{\vec{\vartheta}}-L^{*}\vec{\vartheta}+e^{k}p_{zk}^{\text{osc}}(t),$ (6)
$m_{z}\ddot{\vartheta}_{z}=$ $P_{z}-\alpha_{z}(\dot{\vartheta}_{z}+\omega_{s})-p_{s}-p_{zk}^{\text{osc}}(t),$ (7)
where $L^{*}$ is the effective coupling Laplacian on the synchronized cluster with weights $\kappa_{ij}\cos(\vartheta^{*}_{i}-\vartheta^{*}_{j})$.
Note that $p_{s}$ does not appear in the equation for $S$, since the synchronized cluster steady state accounts for the average power injection from the solitary. We also note that the dependence of $L^{*}$ on $p_{s}$ is typically weak (SM I.B).
For $\vartheta_{z}$ to stay bounded, and for $\omega_{s}$ to be the long-term average of the solitary’s frequency, the long-term average $\vartheta_{z}$ has to be zero. This provides us with the condition $0=P_{z}-\alpha_{z}\omega_{s}-p_{s}$. $p_{s}$ is a measure for the mean coupling of $S$ and $z$, while $\omega_{s}$ stands for their incoherence. Note that this relationship between $\omega_{s}$ and $p_{s}$ can be observed in Fig. 1. In synchrony there is a steady high energy flow. In the natural solitary state, the mean energy flow is zero and the system is effectively decoupled. For intermediate solitary states, there is some intermediate amount of energy.
If we now can express $p_{s}$ as a function of $\omega_{s}$, this equation becomes a self-consistent equation for the frequency of the solitary state:
$0=Z(\omega_{s})\coloneqq P_{z}-\alpha_{z}\omega_{s}-p_{s}(\omega_{s}).$ (8)
Additionally, a potentially stable solitary state requires the correct sign of the change in $Z$ with frequency. If we imagine $\omega_{s}$ as slowly varying, we have $m_{z}\dot{\omega}_{s}=Z(\omega_{s})$. Thus, we expect the sign of the derivative $\partial_{\omega_{s}}Z(\omega_{s})$ to be informative about the stability of the solitary state, at least heuristically.
Both $p_{s}$ and $\omega_{s}$ depend on the trajectory of the solution. To obtain the relationship between them, we will leverage averaging theorems Guckenheimer and Holmes (2002) to obtain an approximate solution for which they can be exactly evaluated. To do so, we introduce a timescale $\varepsilon^{-1}$ that splits the slow, non-rotating and the rotating dynamics. By rescaling the time axis we can eliminate $m$ in Eq. (1), and set $\alpha$, $P$, and $\kappa$ to $\varepsilon$, $\varepsilon P$ and $\varepsilon\kappa$ for a small value $\varepsilon$. We choose $\varepsilon$ such that while the rescaled $\kappa$ remains of order one, the eigenvalues of the rescaled effective Laplacian $L^{*}$ for the eigenmodes localized at $k$ are of order $\varepsilon^{-1}$. As a result, we treat $L^{r}=\varepsilon L^{*}$ as being of order one. Such a scaling can be achieved, if node $k$ has a sufficiently high degree (SM I.A).
After this rescaling, Eq. (6) and Eq. (7) become
ϑ¨=¨italic-$\ddot{\vec{\vartheta}}=$ $\varepsilon\left[-\dot{\vec{\vartheta}}+e^{k}p_{zk}^{\text{osc}}(t)\right]-L^{r}\vec{\vartheta},$ (9)
$\ddot{\vartheta}_{z}=$ $\frac{\varepsilon m}{m_{z}}\left[P_{z}-\frac{\alpha_{z}}{\alpha}\dot{\varphi}_{z}-p_{zk}(t)\right].$ (10)
Equation (9) has the form of the perturbation of the Hamiltonian system $\ddot{\vec{\vartheta}}=-L^{r}\vec{\vartheta}$, where the perturbation is due to $\vartheta_{z}$. The appropriate perturbative approach in this case is averaging Guckenheimer and Holmes (2002), where one should first write the system for the slowly varying amplitudes of the network modes, i.e., the amplitudes of the periodic solutions of the unperturbed system, see Eq. (S48). The amplitude dynamics is then decomposed into modes resonant with $\omega_{c}$, the driving frequency in $p_{zk}^{\text{osc}}$, and non-resonant ones. The contributions of both types of modes are small, allowing us to average the dynamics of $\vec{\vartheta}$. We obtain $\vartheta_{k}(t)$ and insert it into Eq. (10). We then average Eq. (10) and undo the rescaling in this equation to obtain the explicit form of the self-consistent Eq. (8):
$Z(\omega_{s})=P_{z}-\alpha_{z}\omega_{s}-\frac{\kappa_{zk}^{2}}{2}\sum_{\ell=1}^{N-1}\frac{\alpha\omega_{c}\left(v_{k}^{[\ell]}\right)^{2}}{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}}.$ (11)
In Eq. (11), the sum runs over eigenmodes $\ell$ of $L^{*}$, with eigenvalue $\lambda^{[\ell]}$ and eigenvector $v^{[\ell]}$. To efficiently evaluate Eq. (11), we assume that the dependence of $L^{*}$ and its eigenmodes on $p_{s}$ is weak. This is a reasonable assumption if the synchronized cluster is significantly larger than the solitary cluster (node $k$) and the energy injected at node $k$ dissipates quickly into $S$.
A linear response picture.— The obtained form of $p_{s}$ has a natural interpretation. The solitary state, with frequency $\omega_{c}$, perturbs the synchronized cluster. The response of the cluster has the frequency $\omega_{c}$ and some well-defined phase shift. This enables a non-zero energy flow $p_{s}$ between $S$ and $z$, which is responsible for the frequency shift $\omega_{s}$. The flow is proportional to the amplitude of the response and the local amplitude of the corresponding eigenvector. Thus, large energy flows can occur if $\omega_{c}$ is close to resonance with an eigenmode of $S$ that is localized at $k$. As $\omega_{c}$ shifts with the energy flow, the system can robustly tune itself towards such resonances if they are present. This resonant tuning mechanism predicts why certain topological features, namely the ones causing highly localized modes, enable the generation of additional solitary states.
Following this interpretation, we can directly derive the result for $Z(\omega_{s})$ in Eq. (11) using linear response theory, if we neglect the nonlinear part of $p_{zk}^{\text{osc}}$ in Eq. (9), which is eventually averaged out in the averaging approach, from the outset. This alternative assumption can be justified, when this nonlinearity is only one of several contributions to the overall coupling of a well-connected node $k$. In this case, the network is harmonically driven by $\sin\omega_{c}t$, and linear response similar to Zhang et al. (2019) can be used to calculate its response, including the amplitudes $a_{k}^{(k)[\ell]}(\omega_{c})$ and phase lags $\delta^{[\ell]}(\omega_{c})$ at the driven node $k$. These then provide us with a time averaged energy flow:
$p_{s}(\omega_{c})=-\frac{\kappa_{kz}}{2}\sum_{\ell=1}^{N}a_{k}^{(k)[\ell]}(\omega_{c})\sin\delta^{[\ell]}(\omega_{c}).$ (12)
Both approaches used to derive Eq. (11) straightforwardly generalize to a solitary node connected to several nodes in $S$. The back-reactions of the network combine linearly, and we simply sum over all neighbors of $z$ (Eq. (S71)).
Examples.— We illustrate the power of Equation (11) using two examples. First we give a minimal effective model that exhibits the resonant tuning mechanism cleanly; then a full complex network with a complex resonant response. We compare the results with numerical simulations Plietzsch et al. (2022). Further examples can be found in the SM II.B.
Figure 3: Network modes shape the landscape of solitary attractor states for the minimal effective model (inset). (left) Intersections between the straight line $P_{1}-\alpha\omega_{s}$ and the mean energy flow $p_{s}(n=6)$ generate stable and unstable solutions according to Eq. (11). We use $P_{1}=1$, $\alpha=0.1$. The single peak of $p_{s}$ is centered around the network mode close to $\sqrt{Kn}$. (right) Bifurcation diagram with the bifurcation parameter $n$. The two stable branches are connected by an unstable branch and centered around the network mode.
The simplest model that shows tristability with two distinct solitary frequencies is a system that neglects the response of the network beyond the node $k$. We have a solitary node ($z=1$), its neighbor ($k=2$), and the rest of the network is assumed to have infinite inertia, and thus no dynamics. This model is shown in the inset in Fig. 3 (right). The neighbor has $n$ links into the infinitely heavy part, each with coupling strength $K$. See SM II.A for details.
The left panel in Fig. 3 shows the energy absorbed by frequency adaptation, a straight line given by $P_{1}-\alpha\omega_{s}$, and the energy flow due to the network’s response, $p_{s}=\langle p_{12}\rangle$. There is a single peak centered around the only network mode, which is close to $\lambda=\sqrt{Kn}$. At the intersection of these two curves, we have a zero of Eq. (11). However, the linear stability heuristic suggests that only intersections with $\partial_{\omega_{s}}Z<0$ are candidates for stable solitary states. Graphically, these correspond to intersections where the network response intersects from below when reading left to right.
Hence, for $n\lesssim 12$, there is a pair of a stable and an unstable solution due to the peak, and a third, stable solution at $\omega_{s}\approx\Omega_{1}=P_{1}/\alpha$. For larger values of $n$, the second and third solutions annihilate, while the first converges to $\omega_{s}=\Omega_{1}$. The bifurcation diagram in the right panel of Fig. 3 illustrates this behavior. Numerical simulations with random initial conditions closely correspond to the predicted frequency and stability of solitary states, confirming the existence of tristability between the synchronous state and solitary states at multiple frequencies, and the overall accuracy of our results.
We now apply our result to the complex network shown in Figs. 2 and S2. Solitary states with intermediate frequency are located at leaves with neighbor degree $\geq$ 6, (cf. Nitzbon et al. (2017)). Figure 4 shows the solutions of Eq. (11) for $z=74$ for slightly heterogeneous parameters.
Figure 4: Solutions of Eq. (11) for a dense sprout node ($i=74)$ in a synthetic power grid (Fig. 2). For the sake of clarity, we slightly modified the node parameters to reduce the number of intersections in the figure (see SM II.C).
As many network modes are excited, the response curve is considerably more complex. However, simulation results reveal that there is still one dominant mode with high localization at the root node $k$, which is in close resonance with the solitary frequency, and which occurs frequently from random initial conditions.
There are several tentative observations we can make by looking at these examples, (Figs. S4 and S5)): (i) Eq. (11) tends to underestimate the magnitude of the most prominent solution for $\omega_{s}$ slightly, this can already be observed in the effective model; (ii) pairs of close stable and unstable solutions, which are caused by narrow peaks, result in solitary states with a small basin of attraction; (iii) pairs close to their annihilating bifurcation, caused by peaks that barely exceed the straight line at the intersection, also cause small basins; (iv) the most prominent solitary state lies at a frequency that has a high and broad peak and some distance to the next solutions.
Leaf nodes with high neighbor degree Nitzbon et al. (2017), typically produce category (iv) situations (SM I.A). Thus, these findings explain why such nodes feature solitary states with intermediate frequencies.
The derivation of the minimal model, and the explicit relationship between network mode and neighbor degree (SM I.A and II.A), suggests that the neighbor degree can serve as an upper bound of the intermediate frequency in the complex network case. This is also observed numerically (Fig. S8).
Conclusion.— We present an analytical description of the effective coupling between a solitary oscillator and the synchronized cluster in a complex network. To approximate the trajectories, we utilize partial linearization and an averaging theorem. This way, we discover that a resonant excitation of the linear modes of the synchronized cluster can couple coherently to the shifted frequency of the solitary state, resulting in a large energy flow between the synchronized cluster and the solitary oscillator and effectively shifting the frequency.
Our framework predicts the location of solitary oscillations in reduced models, and is in excellent agreement with what we see in complex networks. With this, we can explain that solitary states appear mostly in specific parts of the network Berner et al. (2021), more specifically leaves and tree-shaped structures, which act as weak points in perturbation scenarios Menck et al. (2014); Nitzbon et al. (2017); Halekotte and Feudel (2020); Halekotte et al. (2021). These results have important implications for improving grid stability through modifications of parameters Motter et al. (2013); Wassmer et al. (2021); Menara et al. (2022), topology Witthaut and Timme (2012); Menck et al. (2014); Schultz et al. (2016); Kaiser et al. (2021), and targeted control Taher et al. (2019).
While the presented mechanism can explain the potential existence and stability of intermediate solitary states, there are still open questions that require further investigation. A key practical question is the estimation of the basin boundaries of these states, which would provide deeper insights into the dynamics and robustness of synchronous systems, against perturbations Manik et al. (2014); Klinshov et al. (2015); Halekotte and Feudel (2020); Halekotte et al. (2021); Munyayev et al. (2022).
Detailed numerical studies suggest that additional classes of attractive states exist, and the interaction of losses on the lines and resonant energy flow remains unexplored Hellmann et al. (2020). Furthermore, a more detailed study of the interaction of multiple solitary nodes with distinct frequencies or in frequency clusters is needed.
Finally, while the method introduced above focuses on phase oscillators, it would be interesting to extend it to phase-amplitude oscillators that play a central role in networks such as power grids Kogler et al. (2022). We expect that the overall analytical approach we introduced can be further developed and adapted to address these types of questions.
The code for reproducing our simulations and figures can be found in Niehues (2023). The repository can also be applied to find solitary frequencies in complex networks by solving Eq. (11).
Acknowledgements.
Acknowledgments.— Initial investigations of this phenomenon grew out of work by Anton Plietzsch and discussions with Paul Schultz, Chris Bick, Jeroen Lamb, Deniz Eroglu, Tiago Pereira, and Carsten Grabow during a stay at Nesim Mathematics Village. Part of this work was funded by the DFG grant CoCoHype (DFG KU 837/39-2, J.N., J.K., F.H., M.A.). J.N. gratefully acknowledges support by the Studienstiftung des Deutschen Volkes scholarship foundation and Berlin International Graduate School in Model and Simulation based Research (BIMoS) at Technische Universität Berlin.