### v3

# Resonant Solitary States in Complex Networks

Abstract
Partially synchronized solitary states occur frequently when a synchronized system of networked oscillators with inertia is perturbed locally. Several asymptotic states of different frequencies can coexist at the same node. Here, we reveal the mechanism behind this multistability: additional solitary frequencies arise from the coupling between network modes and the solitary oscillator’s frequency, leading to significant energy transfer. This can cause the solitary node’s frequency to resonate with a Laplacian eigenvalue. We analyze which network structures enable this resonance and explain longstanding numerical observations. Another solitary state that is known in the literature is characterized by the effective decoupling of the synchronized network and the solitary node at the natural frequency. Our framework unifies the description of solitary states near and far from resonance, allowing to predict the behavior of complex networks from their topology.
I Introduction
Many natural and human-made systems are characterized by various degrees of synchronization, which is one of the most fundamental common aspects of their collective behavior Pikovsky et al. (2001). Therefore, important extended systems as diverse as the heart, the brain, firefly populations, chemical reactions, and power grids, are ubiquitously modelled as networks of interconnected oscillators Strogatz (2000); Acebrón et al. (2000, 2005); Pecora et al. (2014); Rodrigues et al. (2016). In some, such as power grids, global synchronization is essential for their proper functioning, in others, such as the brain, it can indicate severe dysfunction. The tendency to synchronize, and thus the function of the system, is strongly influenced by the underlying network’s topology.
The paradigmatic models for understanding this relationship between structure and function in coupled oscillator networks are the Kuramoto model Kuramoto (1984) and its variants. These models feature extremely rich collective behavior, such as chimera states, frequency clusters, isolated desynchronization, and spatial chaos Strogatz (2000); Acebrón et al. (2000, 2005); Pecora et al. (2014); Rodrigues et al. (2016); Munyayev et al. (2023); Kovalenko et al. (2021); Kuramoto (1984). The Kuramoto model with inertia Tanaka et al. (1997a, b) has been developed independently to study synchronization properties in power grids Bergen and Hill (1981); Witthaut et al. (2022) and biological systems Ermentrout (1991).
The transition from decoupling to synchrony in Kuramoto networks is characterized by spatial chaos, a form of extreme multistability Omelchenko et al. (2011). This regime typically has several states with similarly sized basins of attraction Gelbrecht et al. (2020); Halekotte et al. (2021). Among the most prominent are frequency clusters Berner et al. (2021); Olmi et al. (2014), the simplest of which are solitary states. In a solitary state, only one or a few independent oscillators are phase-shifted Maistrenko et al. (2014) or, in the presence of inertia, start to rotate at their own frequencies Jaros et al. (2015), while the rest of the network remains synchronized. Solitary states are especially prominent for the important situation of localized large perturbations Menck et al. (2014); Hellmann et al. (2020); Jaros et al. (2018), and they are particularly likely for perturbations at leaf nodes Menck et al. (2014); Nitzbon et al. (2017). Localized perturbations are important in systems like power grids, where single component failures are common and can lead to desynchronization and blackouts UCTE .
Interestingly, numerical studies have revealed that solitary nodes can exist not only at the natural frequency of the oscillators, but also at intermediate frequencies Nitzbon et al. (2017); Halekotte et al. (2021). In addition, the presence of multistability in these solitary nodes has been identified, allowing for the coexistence of both intermediate and natural solitary states Nitzbon et al. (2017); Halekotte et al. (2021) at the same node.
Previous work has discussed the role of a networks’ topology in understanding its overall response to perturbations and the emergence of various stable states Rohden et al. (2012); Menck et al. (2014); Rodrigues et al. (2016); Nitzbon et al. (2017); Zhang et al. (2019). In some contexts, the frequency and stability of the solitary states have been studied in terms of decoupling and entrainment arguments Menck (2014); Hellmann et al. (2020). These explanations do not account for either intermediate frequency solitaries or the coexistence of several solitary states at the same node.
In this work, we use averaging theory Guckenheimer and Holmes (2002) and linear response to develop a theory of the resonant coupling between the synchronized cluster and the solitary node. We uncover that the described phenomena arise from the cluster’s resonantly excited complex network structure. This leads to a non-zero mean energy flow between the synchronized cluster and the solitary node, and a frequency shift (Fig. 1). While works like Gao and Efstathiou (2018); Yue et al. (2020); Munyayev et al. (2022) have previously considered the interaction between individual oscillators and the mean field, such an interaction between collective network modes and individual oscillators has not been previously described. Here, we uncover how the network’s topology shapes the landscape of solitary states and illustrate our findings with a range of example systems, ranging from conceptual models to fully complex networks.
II Model
Figure 1: Example trajectories of solitary states of intermediate and natural frequencies. Frequencies ($\dot{\varphi}_{i}$) and time-averaged energy flow $p_{s}$ (moving average over 10 time units) between solitary and synchronous component are shown for a complex oscillator network, see Sup for an animation and the network topology (Supplementary Fig. S2). The system is perturbed twice, with the timing indicated by vertical dashed lines. Both perturbations consist of an instantaneous step of $\Delta\dot{\varphi}_{75}=6$ at node 75’s frequency. The first perturbation brings the system from the synchronized regime into an intermediate solitary state in resonance with the synchronized cluster. The second perturbation causes the transition to a natural solitary state, where the solitary node is effectively decoupled and the oscillations of the synchronized cluster are much smaller. The distinct blue trajectory belongs to the solitary node 75 in Fig. 2.
In the following, we will use the Kuramoto model with inertia on complex networks Tanaka et al. (1997a, b); Hellmann et al. (2020); Zhang et al. (2019); Menck et al. (2014); Halekotte and Feudel (2020); Manik et al. (2014). It is given by the coupled second-order equations
$m_{i}\ddot{\varphi}_{i}$ $=P_{i}-\alpha_{i}\dot{\varphi}_{i}-\sum_{j=1}^{N}p_{ij}(\varphi_{i},\varphi_{j}),$ (1)
where $\varphi_{i}$ are the phases of the $N$ oscillators, $P_{i}$ the driving powers, the $m_{i}$ are inertia constants, and the $\alpha_{i}>0$ are damping coefficients. The coupling is given in terms of the weighted coupling matrix $\left\{\kappa_{ij}\right\}$ as
$p_{ij}(\varphi_{i},\varphi_{j})=\kappa_{ij}\sin(\varphi_{i}-\varphi_{j}).$ (2)
Note that the coupling matrix is symmetric, $\kappa_{ij}=\kappa_{ji}$, hence the coupling function is antisymmetric, $p_{ij}=-p_{ji}$. We parametrize the model Eq. (1) such that the synchronous state coexists with various stable attractors of different degree of synchronization. In particular, we set $P_{i}=1$ for producers, $P_{i}=-1$ for consumers, $\alpha_{i}=0.1$, and $\kappa_{ij}=6$ for connected nodes $(i,j)$. This is the parameter regime of, for example, real-world power grids Menck et al. (2014); Nitzbon et al. (2017); Hellmann et al. (2020); Halekotte and Feudel (2020); Halekotte et al. (2021).
To better understand the asymptotic behavior of the system, we recall a few properties of its attractors. In a synchronous state, we have $\varphi_{i}(t)=\varphi_{i}^{*}+\hat{\omega}t$, and constant frequencies $\dot{\varphi}_{i}^{*}=\hat{\omega}$, and the phases satisfy $P_{i}-\alpha_{i}\hat{\omega}=\sum_{j=1}^{N}p_{ij}(\varphi_{i}^{*},\varphi_{j}^{*})$. By going to a corotating frame via $\varphi_{i}\to\varphi_{i}-\hat{\omega}t$, we can always set $\hat{\omega}=0$, and will assume so from now on. The $p_{ij}(\varphi_{i}^{*},\varphi_{j}^{*})$ can be interpreted as the energy flowing through the network to balance out the driving powers $P_{i}$, (this is the physical interpretation for power grids), and $\alpha_{i}$ determines the restoring force of the frequency. The natural frequencies of the uncoupled oscillators (i.e., when all $\kappa_{ij}=0$) are determined by the condition that driving power and restoring force are in balance: $\Omega_{i}\coloneqq P_{i}/\alpha_{i}$. Solitary states in systems of coupled oscillators can be quite similar to the natural solitary states of uncoupled oscillators, which can serve as an approximation Menck et al. (2014).
[…]
To illustrate the phenomenon we describe and explain in this paper, we show simulation results in Fig. 2. It displays the different asymptotic frequencies actually observed following single node perturbations in a complex network, the topology of which was generated with Schultz et al. (2014). The $|\omega_{s}|$, that coincide with $\omega_{s}\cdot\text{sgn}(P_{z})$, are the absolute values of the mean asymptotic solitary frequencies $\omega_{s}$ of the trajectories $\dot{\varphi}_{z}(t)$ at the solitary nodes $z$. The trajectories are won from integrating Eq. (1) from many initial conditions generated by perturbing the synchronous state (see Niehues (2023) for code and details). Here, asymptotic states are counted as 1-solitary if there is exactly one node with frequency $|\omega_{s}|>1$, cf. Fig. S4. Note that many solitary frequencies are close to the natural frequencies at $\Omega_{i}=\pm 10$. We define such solitary states with $\omega_{s}\approx\Omega_{z}$ as natural. However, when perturbations occur at nodes 74, 75 and 84, they can also induce intermediate frequencies. We define solitary states with such intermediate mean frequency $|\omega_{s}|<|\Omega_{z}|$ as resonant, because (i) the fluctuation of the nodes in $S$ is more excited (see Fig. 1), and (ii) because we show that $\omega_{s}$ is in resonance with a network mode. The nodes 74, 75 and 84 share a specific topological property: they are leaf nodes (degree-1 nodes) with a high-degree neighbor. The intermediate resonant solitary states occur at all such nodes in this network (and other networks), and only at these nodes, discarding marginal effects. We explore this strong link between the topology and the asymptotic behavior of the nodes in the following sections.
Figure 2: Local topological features determine the existence and relative occurrence of asymptotic solitary frequencies $\omega_{s}$ in a synthetic power grid observed after perturbations, see Niehues (2023). Three solitary states are clearly distinct from the natural frequencies at ±$\pm 10$; they occur at leaf nodes 74, 75 and 84 which have highly connected neighbors (cf. Nitzbon et al. (2017)) as shown in the insets. Nodes shown as ’$+$’ (’--’) have $P_{i}=1$ $(-1)$. See Fig. S2 for the full network, and Fig. 1 for a solitary state of node 75.
III Ansatz and self-consistency
Here, we present our main result: the framework that allows us to make predictions of solitary states and relate them to the network’s topology. The framework predicts the existence and stability of both natural and resonant solitary states. It consists of an ansatz and a self-consistent equation it has to fulfill.
An ansatz is a proposed form of a solution, often justified by prior knowledge about the system, and its sole purpose is to insert it and evaluate the result it produces. Here, the ansatz is a change of coordinates without loss of generality.
We begin by introducing suitable notation for describing a solitary leaf node $z$ and a synchronized cluster $S=\{i:1\leq i\leq N,i\neq z\}$. For simplicity, we will assume that $\alpha_{i}$ and $m_{i}$ are constant throughout $S$. Node $z$ is connected to $S$ through the intermediary node $k\in S$. With all introduced notations, the considered system has the form
$m\ddot{\varphi}_{i}$ $=P_{i}-\alpha\dot{\varphi}_{i}-\sum_{j=1}^{N}\kappa_{ij}\sin(\varphi_{i}-\varphi_{j}),\quad i\in S,$ (3)
$m_{z}\ddot{\varphi}_{z}$ $=P_{z}-\alpha_{z}\dot{\varphi}_{z}-\kappa_{zk}\sin(\varphi_{z}-\varphi_{k}).$ (4)
The coupling between subsystems $S$ and $z$ is given by $p_{zk}$, which we decompose into a long-term time average $p_{s}\coloneqq\langle p_{zk}\rangle$ and a mean zero part $p_{zk}^{\text{osc}}(t)$:
$p_{zk}(t)=\kappa_{zk}\sin\left[\varphi_{z}(t)-\varphi_{k}(t)\right]\coloneqq p_{s}+p_{zk}^{\text{osc}}(t).$ (5)
Note that we do not have an explicit expression for $p_{s}$ and $p_{zk}^{\text{osc}}(t)$ yet, but obtain them later.
Our ansatz for the trajectories, is to split the linear motion in time with the mean frequency from the mean zero part, similarly to $p_{zk}$ in Eq. (5),
$\varphi_{i}(t)$ $\coloneqq\omega_{\text{sync}}t+\vartheta_{i}^{*}+\vartheta_{i}(t)\quad\text{for $i\in S$,}$ (6)
$\varphi_{z}(t)$ $\coloneqq\omega_{\text{sync}}t+\vartheta_{k}^{*}+\vartheta_{z}(t)\coloneqq\omega_{s}t+\vartheta_{k}^{*}+\varepsilon\psi_{z}(t),$ (7)
for some small $\varepsilon$ that we determine later. We define $\omega_{\text{sync}}$ and $\omega_{s}$ as the long term average frequency of the synchronized cluster, and the solitary oscillator, respectively. This means $\langle\vartheta_{i}(t)\rangle=0=\langle\psi_{z}\rangle(t)$ by construction. The $\vartheta_{i}^{*}$ are a synchronous state of $S$ given a fixed power injection $p_{s}$ at node $k$. Note that the ansatz is without loss of generality, and can be formally considered as a change of variables. We use the form Eqs. (6-7) because if a solitary state occurs, $\psi_{z}(t)$ and $\vartheta_{i}(t)$ are bounded, and the corotating solitary frequency, defined as $\omega_{c}\coloneqq\omega_{s}-\omega_{\text{sync}}$, is non-zero, cf. Supplementary Video 1 Sup . We can use the ansatz Eqs. (5-7) to determine the mean-zero part of the solitary trajectory $\psi_{z}(t)$, however, it is more informative to determine $\omega_{s}$ and $p_{s}$ from the ansatz as a characterization of the solitary state.
To determine $\omega_{s}$, we note that a condition for the existence of a solitary state at $\omega_{s}$ is that both $\omega_{s}$ and $p_{s}$ have to fulfill a self-consistent equation. The argument goes as follows. For $\psi_{z}(t)$ to stay bounded, and for $\omega_{s}$ to be the long-term average of the solitary node’s frequency, the long-term average of $\ddot{\vartheta}_{z}=\varepsilon\ddot{\psi}_{z}$ has to be zero. Demanding this for Eq. (4) provides us with the condition $0=P_{z}-\alpha_{z}\omega_{s}-p_{s}$. We can interpret $p_{s}$ as a measure for the mean coupling of $S$ and $z$, while $\omega_{s}$ stands for their incoherence. Note that this relationship between $\omega_{s}$ and $p_{s}$ can be observed in Fig. 1: in synchrony ($\omega_{s}=0$) there is a steady high energy flow $p_{s}=P_{z}$; in the natural solitary state, the mean energy flow $p_{s}\approx 0$ and the system is effectively decoupled. For intermediate resonant solitary states, there is some intermediate amount of energy transfer $0<|p_{s}|<|P_{z}|$.
If we now can express $p_{s}$ as a function of $\omega_{s}$, the above condition for a solitary state to have well-defined frequency becomes a self-consistent equation for $\omega_{s}$:
$0=Z(\omega_{s})\coloneqq P_{z}-\alpha_{z}\omega_{s}-p_{s}(\omega_{s}).$ (8)
The self-consistency function $Z(\omega_{s})$ is defined as the average change of the solitary frequency as a function of the solitary mean frequency: $Z(\omega_{s})\coloneqq\langle m_{z}\ddot{\vartheta}_{z}\rangle(\omega_{s})$. A potentially stable solitary state at $\omega_{s}$ requires the correct sign of the change in $Z$ with frequency: if a change in $\omega_{s}$ leads to a larger change in $\omega_{s}$ in the same direction, the solitary frequency is linearly unstable, and vice versa. If we imagine $\omega_{s}$ as slowly varying, we have $m_{z}\dot{\omega}_{s}=Z(\omega_{s})$. Thus, we expect the sign of the derivative $\partial_{\omega_{s}}Z(\omega_{s})$ to be informative about the stability of the solitary state, at least as a proxy. In other words, by introducing the explicitly time-dependent ansatz Eqs. (6-7), we get a non-autonomous system from Eqs. (3-4). We are looking for limit cycles of this system that correspond to fixed points in the long-term averaged dynamics. These fixed points are solutions $\omega_{s}$ of Eq. (8): mean frequencies that are allowed to persist by the dynamics, and that are stable under small perturbations.
To obtain the solutions, our strategy is as follows. We assume that we can linearize Eqs. (3-4) in $\vartheta_{i}(t)$ and $\varepsilon$ to obtain a system that quantitatively reflects the behavior of the trajectory of the solitary state. We will analyze the linearized system using averaging, and see that this is justified in many regimes. Together with the self-consistent equation, we obtain a proxy system for the description of solitary states. If the proxy system is stable, we conclude that a solitary state can (but is not guaranteed to) exist at $\omega_{s}$. If the proxy system does not admit a stable solution for specific parameters, we interpret this as evidence that a solitary state can not exist with those parameters. Numerically, we see that if the proxy system admits a stable solution, we often do find solitary states with a large basin. We interpret this to mean that the mechanism revealed by the proxy system is indeed responsible for the formation and stabilization of solitary states.
To follow the strategy outlined above and evaluate Eq. (8) to get predictions for solitary frequencies and their stability, we first need to determine $p_{s}$ as a function of $\omega_{s}$. We present a derivation in the next section, and an interpretation and alternative derivation thereafter.
IV Averaging
Here, we outline the derivation of our main result, an explicit form for $p_{s}(\omega_{s})$. For a detailed step-by-step instruction, we refer to the Supplemental Material Sup , section I (SM I).
Both $p_{s}$ and $\omega_{s}$ depend on the trajectories $\varphi_{z}(t)$ and $\varphi_{k}(t)$ of the solution to Eqs. (3-4). To obtain an explicit approximation of the solution, we will leverage a version of the averaging theorem Guckenheimer and Holmes (2002) (and references therein).
We start by giving a brief summary of the averaging method following Guckenheimer and Holmes (2002). Let $0\leq\varepsilon\ll 1$, and $x\in U\subseteq\mathbb{R}^{n}$ for a bounded set $U$. Let $f:\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{n}$ be a $C^{r}$ function with $r\geq 2$ (at least twice continuously differentiable) that is time-periodic with period $T>0$, and that determines the non-autonomous periodic dynamical system
$\dot{x}=\varepsilon f(x,t,\varepsilon).$ (9)
Define the associated autonomous averaged system as
[…]
The averaging theorem states that a system of the form Eq. (9) can be cast into the form
$\dot{y}=\varepsilon\overline{f}(y)+\varepsilon^{2}f_{1}(y,t,\varepsilon),$ (11)
where $f_{1}$ is also of period $T$ in $t$. This can be achieved by a coordinate change. Moreover, solutions $x(t)$ of Eq. (9) and $y(t)$ of Eq. (10) that are $\varepsilon$-close stay $\varepsilon$-close on a timescale $t\sim\varepsilon^{-1}$. Further, hyperbolic fixed points of Eq. (10) have corresponding unique $\varepsilon$-close hyperbolic periodic orbits of the same stability type. We will leverage this theorem as follows: By using the ansatz Eqs. (6-7) and appropriate approximations, we cast the system Eqs. (3-4) into the form Eq. (9). In the proper coordinate frame, the system is time-periodic and slowly varying. We can then find approximate solutions by solving its corresponding averaged system for fixed points. Those solutions, in particular $\vartheta_{k}(t)$, enable us to calculate $p_{s}$ as a function of $\omega_{s}$ and close the self-consistent equation, Eq. (8).
First, we need to approximate the highly nonlinear dynamics in Eqs. (3-4) appropriately. Since the synchronized oscillators from the cluster $S$ are weakly perturbed by the solitary rotation, we assume smallness of $\vartheta_{i}(t)$, which is also justified by numerical observations, see Fig. 1 and Sup ; Niehues (2023). Therefore, we linearize the system Eqs. (3-4) with respect to $\vartheta_{i}(t)\in S$ around the origin. It is crucial not to linearize with respect to $\vartheta_{z}(t)$, because it is not bounded. For example, the coupling function $p_{zk}$ in Eqs. (4) and (5) is approximated to linear order in $\vartheta_{k}$ as
$p_{zk}(t)$ $\approx\kappa_{zk}\left\{\sin\left[\omega_{c}t+\varepsilon\psi_{z}(t)\right]-\vartheta_{k}(t)\cos\left[\omega_{c}t+\varepsilon\psi_{z}(t)\right]\right\}.$ (12)
From now on, $p_{s}$ and $p_{zk}^{\text{osc}}(t)$ represent the mean and oscillating part of the approximated $p_{zk}$. To write the obtained system in vector form, we define the vectors $e^{k}_{i}:=\delta_{ik}$, where $\delta_{ik}=1$ if $i=k$, and $\delta_{ik}=0$ else. Introducing $\vec{\vartheta}(t)$ with components $\vartheta_{i}(t)$ for all $i\in S$, the linearized system is
$m\ddot{\vec{\vartheta}}=$ $-\alpha\dot{\vec{\vartheta}}-L^{*}\vec{\vartheta}+e^{k}p_{zk}^{\text{osc}}(t),$ (13)
$m_{z}\ddot{\vartheta}_{z}=$ $P_{z}-\alpha_{z}(\dot{\vartheta}_{z}+\omega_{\text{sync}})-p_{s}-p_{zk}^{\text{osc}}(t),$ (14)
where $L^{*}$ is the effective coupling Laplacian on the synchronized cluster with weights $L^{*}_{ij}=\kappa_{ij}\cos(\vartheta^{*}_{i}-\vartheta^{*}_{j})$ and eigenvectors $\vec{v}^{[\ell]}$. We have successfully transformed into a frame that accounts for the power imbalance due to $p_{s}$. This can be seen from the fact that $p_{s}$ does not appear in the equation for $S$, Eq. (13), since the synchronized cluster steady state accounts for the average power injection $p_{s}$ from the solitary node with the $\omega_{\text{sync}}$ rotation.
Second, we need to properly take care of the fact that Eqs. (13-14) become a non-autonomous system when we insert the explicitly time-dependent ansatz Eqs. (6-7) into $p_{zk}$. We introduce a timescale $\varepsilon^{-1}$ that splits the autonomous and non-autonomous parts in Eqs. (13-14). It allows us to write the system in a perturbative form, in which averaging can be applied. Such a scaling can be achieved, if node $k$ has a sufficiently high degree (SM I.A). In that case, there is a Laplacian eigenvalue $\lambda^{[r]}$ that is highly correlated to the node’s degree $d_{k}$. We will observe in the examples below that the solitary frequency tunes itself towards resonance with such eigenmodes: $\omega_{s}^{2}\sim\lambda^{[r]}$.
The rescaling is given by
$t=\sqrt{\frac{m}{\lambda^{[r]}}}t^{\prime},\qquad P_{i}=\frac{P^{\prime}_{i}\alpha\sqrt{\lambda^{[r]}}}{\sqrt{m}},\qquad\kappa_{ij}=\frac{\kappa^{\prime}_{ij}\alpha\sqrt{\lambda^{[r]}}}{\sqrt{m}},$ (15)
dropping the prime from now on. Further, we set $\varepsilon=\alpha/\sqrt{m\lambda^{[r]}}\ll 1$. In other words, by rescaling the time axis we can absorb $m$ in Eq. (13), and set $\alpha$, $P$, and $\kappa$ to $\varepsilon$, $\varepsilon P$ and $\varepsilon\kappa$. We choose $\varepsilon$ such that, while the rescaled $\kappa$ remains of order one, the eigenvalues of the rescaled effective Laplacian $L^{*}$ for the eigenmodes localized at $k$ are of order $\varepsilon^{-1}$. As a result, we treat $L^{r}=\varepsilon L^{*}$ as being of order one. Eqs. (13-14) become
ϑ¨=¨italic-$\ddot{\vec{\vartheta}}=$ $\varepsilon\left[-\dot{\vec{\vartheta}}+e^{k}p_{zk}^{\text{osc}}(t)\right]-L^{r}\vec{\vartheta},$ (16)
$\ddot{\vartheta}_{z}=$ $\frac{\varepsilon m}{m_{z}}\left[P_{z}-\frac{\alpha_{z}}{\alpha}\dot{\varphi}_{z}-p_{zk}(t)\right].$ (17)
Third, a crucial step for solvability is to approximate $\varepsilon p_{zk}(t)$, cf. Eq. (12), to first order in $\varepsilon$, which finally allows us to identify an explicit expression for $p_{s}$ in terms of $\vartheta_{k}$ and $\omega_{c}$,
$\varepsilon p_{zk}(t)$ $\approx\varepsilon\kappa_{zk}\left(\sin\omega_{c}t-\vartheta_{k}(t)\cos\omega_{c}t\right),$ (18)
$p_{s}$ $\approx\kappa_{zk}\langle\vartheta_{k}(t)\sin\omega_{c}t\rangle.$ (19)
Equation (16) has the form of the perturbation of the Hamiltonian system $\ddot{\vec{\vartheta}}=-L^{r}\vec{\vartheta}$, where the perturbation is of order $\varepsilon$. The perturbation is caused partly by the presence of $\vartheta_{z}(t)$, and by other small terms, such as damping. The appropriate perturbative approach in this case is averaging Guckenheimer and Holmes (2002), where one should first write the system for the slowly varying amplitudes $\vec{x}(t)$ and $\vec{z}(t)$ of the network modes $\vec{\xi}(t)$, i.e., the amplitudes of the periodic solutions of the unperturbed system. This is achieved by diagonalization of the Laplacian with the linear transformation $Q_{ij}\coloneqq v_{i}^{[j]}$,
$\vec{\xi}(t)\coloneqq Q^{T}\vec{\vartheta}(t)\coloneqq\vec{x}(t)\cos\omega_{c}t-\vec{z}(t)\sin\omega_{c}t.$ (20)
The amplitude dynamics for $\vec{x}(t)$ and $\vec{z}(t)$ can be inferred with the invertible van der Pol transformation (spelled out in Eq. (S50)) into the frame rotating at the driving frequency $\omega_{c}$. These amplitude dynamics is then decomposed into modes resonant with $\omega_{c}$ and non-resonant modes. For the sake of brevity and clarity, we refer to SM I for details. Most importantly, the contributions of both types of modes are small for different reasons. On one hand, the resonant modes give small contributions due to a small prefactor $\varepsilon\Delta\sim\omega_{c}^{2}-\lambda^{[r]}$. On the other hand, the non-resonant modes, where $\varepsilon\Delta$ is of $O(1)$, give small contributions due to their weak localization, allowing us to average the dynamics of the slowly varying $\vec{x}(t)$ and $\vec{z}(t)$. We obtain an autonomous system that is a good approximation ($\varepsilon$-close) of the non-autonomous system for time scales up to $\varepsilon^{-1}$ Guckenheimer and Holmes (2002). This averaged system (Eq. (S55)) is linear with a unique solution that gives us the average amplitudes $\vec{x}^{*}$ and $\vec{z}^{*}$ of all modes. From those, we obtain the average trajectory of $\vartheta_{k}(t)$, and insert it into Eq. (19) to get an expression for $p_{s}(\omega_{s})$. This finally provides us with the explicit form of the self-consistent Eq. (8):
$0=Z(\omega_{s})$ $=P_{z}-\alpha_{z}\omega_{s}-p_{s}$ (21)
$=P_{z}-\alpha_{z}\omega_{s}-\frac{\kappa_{zk}^{2}}{2}\sum_{\ell=1}^{N-1}\frac{\alpha\omega_{c}\left(v_{k}^{[\ell]}\right)^{2}}{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}}.$
In Eq. (21), the sum runs over eigenmodes $\ell$ of $L^{*}$, with eigenvalue $\lambda^{[\ell]}$ and eigenvector $\vec{v}^{[\ell]}$. To efficiently evaluate Eq. (21), we assume that the dependence of $L^{*}$ and its eigenmodes on $p_{s}$ is weak. This is a reasonable assumption if the synchronized cluster is significantly larger than the solitary cluster (node $z$), and the energy injected at node $k$ dissipates quickly into $S$. Practically, the $\vartheta_{i}^{*}$, that depend on $p_{s}$, can be well approximated by the $\varphi_{i}^{*}$, that do not (SM I.B).
Lastly, to close Eq. (21), we need to relate the solitary frequency $\omega_{s}$ to $\omega_{c}$, its relative value to $S$. It can be shown by summing over Eq. (1) for all $i\in S$ and $i=z$, that the $\alpha$-weighted sum of frequencies decays towards zero from all initial conditions, hence $(N-1)\alpha\omega_{\text{sync}}+\alpha_{z}\omega_{s}=0$ (see SM I.A.). This gives us
$\omega_{c}(\omega_{s})=\left(1+\frac{\alpha_{z}}{(N-1)\alpha}\right)\omega_{s}.$ (22)
This linear relationship between $\omega_{s}$ and $\omega_{c}$ is useful in solving Eq. (21).
Having obtained an explicit, closed, self-consistent equation for $\omega_{s}$ from our ansatz, we now turn to its interpretation and evaluation to make predictions of solitary attractor states and compare them to simulations.
V Interpretation: a linear response picture
In this section, we provide an alternative derivation that is a shortcut to Eq. (21), using linear response theory. For an extensive account of linear response in inertial Kuramoto networks, we refer to Zhang et al. (2019). As a motivation, we start by discussing the natural interpretation that the linear response offers, and that can also be deducted from the resulting Eq. (21).
The obtained form of $p_{s}$ in Eq. (21) has a natural interpretation. The solitary state, with frequency $\omega_{s}$, perturbs the synchronized cluster $S$, with relative frequency $\omega_{c}$. The response $\vec{\vartheta}(t)$ of the cluster $S$ oscillates at frequency $\omega_{c}$ with amplitudes $a_{i}$ and some well-defined phase shifts $\delta_{i}$, e.g., $\vartheta_{k}(t)\sim a_{k}\sin\left(\omega_{c}t+\delta_{k}\right)$. The phase shift $\delta_{k}$ enables a non-zero energy flow $p_{s}$ between $S$ and $z$, which is responsible for maintaining the relative frequency shift $\omega_{c}$. The energy flow is proportional to the amplitude of the response, and especially the local amplitude of the corresponding eigenvector $v_{k}^{[\ell]}$ for each mode $\ell$. Thus, large energy flows can occur if $\omega_{c}$ is close to resonance with an eigenmode of $S$ that is localized at $k$. As $\omega_{c}$ determines the magnitude of energy flow, but also shifts with the energy flow itself, the system can robustly tune itself towards such resonances if they are present. This resonant tuning mechanism predicts why certain topological features, namely the ones causing highly localized modes, enable the generation of additional solitary states, cf. section VI.
Following this interpretation, we can directly derive the result for $Z(\omega_{s})$ in Eq. (21) using linear response theory, see SM I.C. We observe that the nonlinear part of $p_{zk}^{\text{osc}}(t)$ in Eq. (16) is eventually averaged out in the averaging approach. Linear response can be readily applied, if we neglect this nonlinearity from the outset. The assumption that this nonlinearity is sufficiently small can be justified, when it is only one of several contributions to the overall coupling of a well-connected node $k$. In this case, the cluster $S$ is harmonically driven by $\sin\omega_{c}t$, and linear response similar to Zhang et al. (2019) can be used to calculate its response for each network mode $\ell$ separately, including the modal amplitudes $a_{i}^{(k)[\ell]}(\omega_{c})$ at node $i$ and modal phase lags $\delta^{[\ell]}(\omega_{c})$ given by
$a_{i}^{(k)[\ell]}(\omega_{c})$$\coloneqq$ $\frac{\kappa_{kz}v_{i}^{[\ell]}v_{k}^{[\ell]}}{\sqrt{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}}},$ (23)
$\sin\delta^{[\ell]}(\omega_{c})$$\coloneqq$ $-\frac{\alpha\omega_{c}}{\sqrt{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}}}.$ (24)
Here, the superscript (k) indicates that the perturbation caused by $\varphi_{z}$ enters $S$ at node $k$. For each mode $\ell$, the nodal amplitudes depend on the mismatch between the driving frequency $\omega_{c}$ and the eigenmodes of the network associated with the $\lambda^{[\ell]}$, as well as the localization of the corresponding eigenvectors at nodes $i$ and $k$. The damping slightly shifts the location of the resonance peaks in the amplitude away from $\omega_{c}=\pm\sqrt{\lambda^{[\ell]}/m}$ and makes their height finite.
Combining the modal responses linearly provides us with a trajectory for the $\vartheta_{i}$,
$\kappa_{kz}\vartheta_{i}^{(k)}(t)=\sum_{\ell=1}^{N-1}a_{i}^{(k)[\ell]}\sin\left(\omega_{c}t+\delta^{[\ell]}\right).$ (25)
From the trajectory of $\vartheta_{k}$ we can determine the time averaged energy flow $p_{s}=\langle p_{zk}\rangle$ in terms of the amplitudes and phase lags:
$p_{s}(\omega_{c})$ $=$ $-\frac{\kappa_{kz}}{2}\sum_{\ell=1}^{N}a_{k}^{(k)[\ell]}(\omega_{c})\sin\delta^{[\ell]}(\omega_{c})$ (26)
$=$ $\frac{\kappa_{zk}^{2}}{2}\sum_{\ell=1}^{N-1}\frac{\alpha\omega_{c}\left(v_{k}^{[\ell]}\right)^{2}}{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}},$ (27)
which is the exact same as in Eq. (21). In fact, the terms in the denominators in the sum over all modes in Eq. (21) resemble the resonance curve of linear oscillators. It is known as Cauchy–Lorentz distribution, Lorentz(ian) function, or Breit–Wigner distribution. Due to the generalization to the resonance of a network $S$, the Lorentzians are summed over all $N-1$ network modes associated to the $\lambda^{[\ell]}$. Each contribution is weighted by the localization of the corresponding eigenvector $\vec{v}^{[\ell]}$ at node $k$, quantified by $\left(v_{k}^{[\ell]}\right)^{2}$. This is how the topology affects the distribution of resonance in the network, see also Zhang et al. (2019).
We remark that both approaches used to derive Eq. (21) straightforwardly generalize to a solitary node connected to several nodes in $S$, see SM I.D. The back-reactions of the network combine linearly, and we simply sum over all neighbors of $z$ (Eq. (S74)).
Furthermore, a treatment of heterogeneous system parameters within $S$ is also possible, see SM I.E.
VI Evaluation
Here, we make some general statements about the properties and solutions of our main result, Eq. (21), and provide some examples.
First, note that the obtained expression for $p_{s}=\langle p_{zk}\rangle$ in Eq. (21), has the same sign as $\omega_{c}$. Therefore, $p_{s}$ is antisymmetric in $\omega_{c}$ and hence in $\omega_{s}\sim\omega_{c}$, cf. Eq. (22). It follows that Eq. (21) is invariant under $\omega_{s}\to-\omega_{s}$ while $P_{z}\to-P_{z}$, i.e., generators and consumers behave similarly with their solitary frequencies having opposite signs. In fact, it is easy to see from Eq. (21) that all solutions lie between zero and the natural frequency, $0\leq|\omega_{s}|\leq|\Omega_{z}|=|P_{z}|/\alpha_{z}$. We see these bounds and symmetry properties confirmed in Fig. 2, and the examples below.
As discussed above, the linear stability of solutions of $Z(\omega_{s})=0$ can be heuristically determined by evaluating the derivative: the orbit at $\omega_{s}$ is indicated to be linearly stable if $\partial_{\omega_{s}}Z(\omega_{s})<0$ and linearly unstable if $\partial_{\omega_{s}}Z(\omega_{s})>0$. This consideration gains importance with the numerical observation that the instantaneous solitary frequency $\dot{\varphi}_{z}(t)$ oscillates with time around $\omega_{s}$, cf. Fig. 1. The solution to $Z(\omega_{s})=0$ is an intersection of $p_{s}(\omega_{s})$ and the straight line $P_{z}-\alpha_{z}\omega_{s}$ with slope $-\alpha_{z}$. Therefore, stability is given if $-\alpha_{z}<\partial_{\omega_{s}}p_{s}(\omega_{s})$. Graphically, this means that $p_{s}(\omega_{s})$ intersects the straight line from below, counting in positive $\omega_{s}$ direction.
The existence of solitary frequencies $\omega_{s}$ as solutions to Eq. (21) relies on sufficiently high eigenvector localization at the neighboring node $k$, quantified by $\left(v_{k}^{[\ell]}\right)^{2}$. Such high localization is present in random networks of high degree heterogeneity Hata and Nakao (2017). Indeed, we observe in Fig. 2 and in the examples below that resonant solitary states are most common at degree-one nodes with a high degree neighbor Nitzbon et al. (2017), and that the solitary frequency scales with the degree $d_{k}$ of node $k$. For an estimate of the relation between the eigenvector localization, the corresponding eigenvalue and the node’s degree, we refer to SM I.A.
The existence of solutions to Eq. (21) also depends on system parameters and their topological distribution. Whether and which solitary states exist in a complex system is a complex question with no easy answer, but we can provide some insights (see SM II.D for more details). Generally, the parameters need to be chosen in an intermediate regime, such that the synchronous state is not globally attractive, but the coupling is also not too weak. Relatively small damping $\alpha_{i}$ and strong coupling $\kappa_{ij}$, such as in power grids, are an indicator for this. We remark that solving the self-consistent Eq. (21) is an efficient way to determine which solitary states exist. Numerical simulations of the dynamics as in Fig. 2 are more precise but more expensive.
We now illustrate the power of Eq. (21) using two examples. First we give a minimal effective model that exhibits the resonant tuning mechanism cleanly; then a full complex network with a complex resonant response. We compare the results with numerical simulations Plietzsch et al. (2022). Further examples can be found in the SM II.B.
Figure 3: Network modes shape the landscape of solitary attractor states for the two-node model (inset). (left) Intersections between the straight line $P_{1}-\alpha\omega_{s}$ and the mean energy flow $p_{s}(n=6)$ generate stable and unstable solutions according to Eq. (21). We use $P_{1}=1$, $\alpha=0.1$. The single peak of $p_{s}$ is centered around the network mode close to $\sqrt{Kn}$. (right) Bifurcation diagram with the bifurcation parameter $n$. The two stable branches are connected by an unstable branch and centered around the network mode. Numerical simulations Niehues (2023) confirm predictions of both the location and stability of solitary states.
VI.1 Example I: Two-node model
A minimal model that still shows tristability with two distinct solitary frequencies is a system that neglects the response of the network beyond the node $k$. We have a solitary node ($z=1$), its neighbor ($k=2$), and the rest of the network is assumed to have infinite inertia, and thus no dynamics. This model is shown in the inset in Fig. 3 (right). The neighbor has $n$ links into the infinitely inert part, each with coupling strength $K$. See SM II.A for details.
The left panel in Fig. 3 shows the graphical solution of Eq. (21) for this example system. We see the energy absorbed in node 1 by frequency adaptation, a straight line given by $P_{1}-\alpha\omega_{s}$, and the energy flow due to the network’s response, $p_{s}(\omega_{s})=\langle p_{12}\rangle(\omega_{s})$. The energy flow has a single peak centered around the only network mode, which is close to $\lambda\approx Kn$. At the intersections of these two curves, we have zeros of Eq. (21). However, the linear stability heuristic suggests that only intersections with $\partial_{\omega_{s}}Z<0$ are candidates for stable solitary states. Indeed, numerical simulations confirm the existence of the stable solutions, and that they correspond to the intersections, where $p_{s}$ comes from below.
Repeating this solution process for Eq. (21) for a range of $n$, we can draw the bifurcation diagram in Fig. 3 (right). Each vertical slice corresponds to the top view of an intersection plot like the one in the left panel. For this minimal example system, we can obtain an explicit expression for several branches $n(\omega_{s})$ of stable or unstable solutions (see SM II.A) that correspond to the intersections in the left panel. These branches predict that for $n\lesssim 12$, there is a pair of a stable and an unstable solution $\omega_{s}$ due to the resonance peak around $\lambda$, and a third, stable solution at $\omega_{s}\approx\Omega_{1}=P_{1}/\alpha$. For larger values of $n$, the second and third solutions annihilate, while the first converges to $\omega_{s}\to\Omega_{1}$. Numerical simulations with random initial conditions Niehues (2023) closely correspond to the predicted frequency and stability of solitary states, confirming the existence of tristability between the synchronous state and solitary states at multiple frequencies, and confirming the overall accuracy of our results.
In summary, this minimal example system shows that the resonant (intermediate) solitary state and the natural (decoupled) solitary state are two ends of a spectrum of frequencies, and connected by an unstable branch of solutions. Further, the natural frequency is an upper bound for solitary frequencies and can be approached by tuning the resonant network mode higher.
VI.2 Example II: Complex network
We now apply our result to the complex network shown in Figs. 2 and S2. For models of this complexity, the self-consistent equation has to be solved numerically for every node. We observe that intermediate resonant solitary states are located at leaf nodes with neighbor degree $\geq$ 6, (cf. Nitzbon et al. (2017)). Figure 4 shows the solutions of Eq. (21) for $z=74$ for slightly heterogeneous parameters.
Figure 4: Solutions of Eq. (21) for a dense sprout node ($i=74)$ in a synthetic power grid (Fig. 2). For the sake of clarity, we slightly modified the node parameters to reduce the number of intersections in the figure (see SM II.C).
As many network modes are excited, the response curve $p_{s}(\omega_{s})$ is considerably more complex than for the two-node model. However, simulation results reveal that there is still one dominant mode with high localization at the root node $k$. Solitary states obtained from random initial conditions are frequently found in close resonance with the dominant mode, or close to the natural frequency.
There are several tentative observations we can make by looking at complex networks, (see also SM II.C, Figs. S5 and S6 for more examples): (i) If the self-consistent equation yields more solitary states than the one at the natural frequency, there are two dominating solitary states: the one at the natural frequency, and one in resonance with a localized network mode. There can also be many additional solutions with small basins of attraction, cf. Figs. S4 and S5.
(ii) Eq. (21) tends to overestimate the magnitude of the most prominent solution for $\omega_{s}$ slightly, which can already be observed in the effective model (cf. Figs. 3 and 4). (ii) the most prominent natural solitary state lies at an intermediate frequency that has a high and broad peak in $p_{s}$ and some distance to the next solutions. This can be understood by considering that every peak in $p_{s}$ produces a pair of solutions, a stable one and an unstable one, and that the solitary frequency $\varphi_{z}(t)$ is modulated around $\omega_{s}$, cf. Fig. 1. Leaf nodes with high neighbor degree Nitzbon et al. (2017), typically produce such situations (cf. SM I.A) with a highly localized network mode that results in a high peak. Thus, these findings explain why such nodes feature resonant solitary states with intermediate frequencies. (iii) Besides the main peak at the most localized network mode, there can be several minor peaks that result in solitary states with much smaller basins of attraction due to three reasons: First, the peaks might be narrow, such that the stable and unstable solution are close, second, the peak might not exceed the intersection with the straight line by much, hence the pair of solutions is already close to their annihilation bifurcation, and third, the peak might be close to other peaks and their solutions. This can account for the observation that parts of the phase space can be dominated by basin boundaries of many, different, unlikely states Gelbrecht et al. (2020), inducing transient chaos Halekotte et al. (2021).
As a closing remark, the derivation of the minimal model, and the explicit relationship between network mode and neighbor degree (SM I.A and II.A), suggest the following. Even in complex networks, the neighbor degree $d_{k}$ can serve as an upper bound for the (squared) intermediate solitary frequency, when weighted with the coupling $\kappa_{ij}$. This is also observed numerically (Fig. S8).
VII Summary and discussion
We present an analytical description of the effective coupling between a solitary oscillator and the synchronized cluster in a complex network. To approximate the trajectories, we utilize partial linearization and an averaging theorem. This way, we discover that a resonant excitation of the linear modes of the synchronized cluster can couple coherently to the shifted frequency of the solitary state, resulting in a large energy flow between the synchronized cluster and the solitary oscillator, and effectively shifting the frequency.
We have mainly found that intermediate solitary states, first observed in Nitzbon et al. (2017), are actually resonant solitary states, and we have uncovered how the network’s topology shapes their properties. Furthermore, we have unified their analytical description with the one of natural solitary states Menck et al. (2014), and shown that there is a spectrum connecting those extremes.
Our self-consistent framework predicts solitary states in reduced models, where it is solvable, and is also in excellent agreement with what we observe numerically in the original complex networks. With this structure-function relationship, we can explain the observation that solitary states appear mostly in specific parts of the network Berner et al. (2021), more specifically leaf nodes and tree-shaped structures, which act as weak points in perturbation scenarios Menck et al. (2014); Nitzbon et al. (2017); Halekotte and Feudel (2020); Halekotte et al. (2021). These results have important implications for improving grid stability through modifications of parameters Motter et al. (2013); Wassmer et al. (2021); Menara et al. (2022) and topology Witthaut and Timme (2012); Menck et al. (2014); Schultz et al. (2016); Kaiser et al. (2021): Our framework can be used to evaluate how these modifications affect the presence and stability of potentially harmful solitary attractors and their proximity to the synchronous state. Furthermore, we can efficiently identify the troublemaker nodes that tend to desynchronize the easiest due to their intermediate frequencies: resonant solitary states at certain leaf nodes (dense sprout nodes, see Sup ; Nitzbon et al. (2017)). We can identify them from the network topology only, and without expensive simulations. It has been shown that targeted control of such troublemaker nodes can improve grid stability drastically Taher et al. (2019).
VIII Conclusion and Outlook
While the presented mechanism can explain the potential existence and stability of resonant solitary states, there are still open questions that require further investigation. A key practical question is the estimation of the basin boundaries of these states, which would provide deeper insights into the dynamics and robustness of synchronous systems against perturbations Manik et al. (2014); Klinshov et al. (2015); Halekotte and Feudel (2020); Halekotte et al. (2021); Munyayev et al. (2022). It is known that the basin boundaries depend on the phase and frequency Menck et al. (2014); Hellmann et al. (2020); Halekotte et al. (2021). Averaging the phase dependence out, an estimate of the basin boundaries in $\omega$-direction could be estimated by the turning points of $Z(\omega)$ around a stable solution.
Detailed numerical studies suggest that additional classes of attractive states exist, and the interaction of losses on the lines and resonant energy flow remains unexplored Hellmann et al. (2020). Furthermore, a more detailed study of the interaction of multiple solitary nodes with distinct frequencies or in frequency clusters is needed (see SM I.F for an outline).
Finally, while this paper focuses on explaining a known phenomenon of Kuramoto oscillators, which exhibit a purely sinusoidal coupling, our derivation generally only demands a $2\pi$-periodic coupling function. This includes coupling functions with phase shifts and harmonics. It would also be interesting to treat phase-amplitude oscillators that play a central role in networks such as power grids Kogler et al. (2022). We expect that the overall analytical approach we introduced is a promising tool in describing the synchronization of rotation and oscillation, and leave it to future work to develop and adapt our theory to address a wide range of models.
Code Availability
The numerical experiments were conducted with the PowerDynamics.jl package Plietzsch et al. (2022). Further details can be found in SM III. The code for reproducing our simulations and figures can be found in Niehues (2023). The repository can also be applied to find solitary frequencies in complex networks by solving Eq. (21).
Acknowledgements.
Acknowledgments
Initial investigations of this phenomenon grew out of work by Anton Plietzsch and discussions with Paul Schultz, Chris Bick, Jeroen Lamb, Deniz Eroglu, Tiago Pereira, and Carsten Grabow during a stay at Nesim Mathematics Village.
Part of this work was funded by the DFG grant CoCoHype (DFG KU 837/39-2; J.N., J.K., F.H., M.A.) and BMWK grant OpPoDyn (03EI1071A; J.N., F.H.).
J.N. gratefully acknowledges support by Studienstiftung des Deutschen Volkes scholarship foundation and Berlin International Graduate School in Model and Simulation based Research (BIMoS) at Technische Universität Berlin. S.Y. and J.K. acknowledge funding by DFG, German Research Foundation, Project No. 411803875.